OpenStudy (anonymous):

Multiply (2x+5)(x-3)

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@Koikkara

OpenStudy (koikkara):

Hey, 2x(x-3)+5(x-3)= ??

OpenStudy (anonymous):

Huh? You want me to solve that?

OpenStudy (koikkara):

yeah, not me...your question says to solve it... ! \(Lol\)

OpenStudy (anonymous):

2x(x-3)+5(x-3)= 2x^2-3+5x-15?

OpenStudy (koikkara):

oh sorry, typed little wrong... 2x^2-x-15=0

OpenStudy (anonymous):

Thanks. :)

OpenStudy (koikkara):

\(Welcome~!!\)

OpenStudy (koikkara):

 \(Reference\)